OpenStudy (jazzyluvsu123):

5/6x=20 How do i divide this the correct way? or do I not divide?

OpenStudy (dinamix):

what u mean @JazzyLuvsU123

ILovePuppiesLol (ilovepuppieslol):

u have to isolate the variable

ILovePuppiesLol (ilovepuppieslol):

multiply 6/5 to both sides, so that the 5/6 on the x side is gone

ILovePuppiesLol (ilovepuppieslol):

and then u will have x= 20 times 6/5 and then solve

OpenStudy (dinamix):

x= \[x=\frac{ 6*20 }{ 5 }\]

ILovePuppiesLol (ilovepuppieslol):

thats one way to do it

OpenStudy (dinamix):

@JazzyLuvsU123

OpenStudy (anonymous):

1/24 is ur answer

OpenStudy (anonymous):

Write the equation. \[\frac{ 5 }{ 6 }x=20\] Multiply both sides by the reciprocal of the coefficient to isolate the variable (x). \[\left( \frac{ 6 }{ 5 } \right)\frac{ 5 }{ 6 }x=\left( \frac{ 6 }{ 5 } \right)\frac{ 20 }{ 1 }\] \[x=\frac{ 120 }{ 5 }\] Simplify. \[x=24\] Tada.